# Decoupling Local Primordial non-Gaussianity from Relativistic Effects in the Galaxy Bispectrum

### v2

@@ Abstract
Upcoming galaxy surveys aim to map the Universe with unprecedented precision, depth and sky coverage. The galaxy bispectrum is a prime source of information as it allows us to probe primordial non-Gaussianity (PNG), a key factor in differentiating various models of inflation. On the scales where local PNG is strongest, Doppler and other relativistic effects become important and need to be included. We investigate the detectability of relativistic and local PNG contributions in the galaxy bispectrum. We compute the signal-to-noise ratio for the detection of the bispectrum including such effects. Furthermore, we perform information matrix forecasts on the local PNG parameter, $f_{\rm NL}$, and on the parametrised amplitudes of the relativistic corrections. Finally, we quantify the bias on the measurement of $f_{\rm NL}$ that arises from neglecting relativistic effects. Our results show that detections of both first- and second-order relativistic effects are promising with forthcoming spectroscopic survey specifications – and are largely unaffected by the uncertainty in $f_{\rm NL}$. Conversely, we show for the first time that neglecting relativistic corrections in the galaxy bispectrum can lead to a shift $>\!1.5\,\sigma(f_{\rm NL})$ on the detected value of $f_{\rm NL}$, highlighting the importance of including relativistic effects in our modelling.
@@ 1 Introduction
Galaxy surveys have long been an indispensable probe of large-scale structure. We can infer the underlying physics that governs the evolution of matter perturbations by observing the galaxy distributions at different scales and redshifts, and therefore test the standard model of cosmology as well as theories of gravity. The anticipated measurements from current and future surveys such as the Dark Energy Spectroscopic Instrument (DESI) [1, 2], the Euclid satellite mission [3, 4], the Nancy Grace Roman Space Telescope [5, 6] and MegaMapper [7], will provide information on ultra-large scales, i.e. beyond the matter-radiation equality scale. At these scales primordial non-Gaussianity (PNG) can be probed. It is also at these scales that effects from general relativity have the most impact on our measurements. This is the key motivation of this paper, quantifying how corrections to the galaxy number density which appear due to general relativity, specifically those derived from observing on our past lightcone, affect measurements of PNG. We explore this through the galaxy bispectrum, the Fourier transform of the 3-point correlation function of galaxy density contrasts, which is only non-zero in the presence of non-Gaussianity. Doppler shifts induced by the peculiar velocities of galaxies, gravitational lensing effects, and other relativistic corrections alter the clustering patterns of galaxies, imprinting distinct signatures on the bispectrum which can mimic those of PNG [8, 9, 10, 11]. Therefore, understanding and accurately modeling these relativistic effects are imperative for extracting robust cosmological information from galaxy surveys.
@@
The galaxy power spectrum has been extensively studied and used to obtain impressive constraints on cosmological parameters. The dominant perturbation effect on sub-Hubble scales is from redshift-space distortions (RSD) [16], which constitute the standard Newtonian approximation to projection effects. The inclusion of relativistic corrections to the galaxy density contrast at linear order [17, 18, 19] has, in turn, provided further insights. Many studies have been done on the possibility of constraining the local PNG parameter $f_{\mathrm{NL}}$ using the relativistic galaxy power spectrum for future galaxy and intensity mapping surveys, which avoids the bias typical in the Newtonian analysis [8, 9, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. PNG in the galaxy bispectrum has been extensively investigated in the Newtonian approximation [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. In [43] it was shown that constraints on $f_{\mathrm{NL}}$ can be obtained from large-scale structure which are competitive with that of Planck, using forecasts on the Newtonian tree-level bispectrum (including RSD to second order), based on survey specifications from BOSS [54], eBOSS [55], Euclid [56] and DESI (LRGs, ELGs and QSOs) [57].
@@ 2 The galaxy bispectrum
$2\,\left<\varDelta_{\rm g}^{(1)}(\bm{k}_{1})\,\varDelta_{\rm g}^{(1)}(\bm{k}_{2})\,\varDelta_{\rm g}^{(2)}(\bm{k}_{3})\right>+2\,\mathrm{cp}=(2\pi)^{3}\,B_{\rm g}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,\delta^{D}(\bm{k}_{1}+\bm{k}_{2}\,+\bm{k}_{3})\;,$ (2.1)
where we use the convention that the observed density contrast is $\varDelta_{\rm g}=\varDelta_{\rm g}^{(1)}+\varDelta_{\rm g}^{(2)}/2$
and is related at first order to the number density contrast at the source $\delta_{\rm g}$ by
$\varDelta_{\rm g}=\delta_{\rm g}+\mathrm{RSD}+\text{relativistic effects}+\text{local PNG contributions}\;.$ (2.2)
It can then be written in terms of first- and second-order Fourier kernels as
$B_{\rm g}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})=\mathcal{K}^{(1)}(\bm{k}_{1})\,\mathcal{K}^{(1)}(\bm{k}_{2})\,\mathcal{K}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{cp}\;,$ (2.3)
@@
$B_{\rm gN}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})=\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{2})\,\mathcal{K}_{\rm N}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{cp}\;,$ (2.4)
@@
Here $\mu_{a}=\hat{\bm{k}}_{a}\cdot\bm{n}$, with $\bm{n}$ the line-of-sight direction, $f=\mathrm{d}\ln D/\mathrm{d}\ln a$ is the linear matter growth rate, and $D$ is the linear growth factor. The kernels $F_{2}$ (mode-coupling part of the matter density contrast), $G_{2}$ (mode-coupling part of the matter velocity), and $S_{2}$ (tidal bias) are given in Section A.1, with the linear and quadratic Gaussian clustering biases $b_{10}$ and $b_{20}$ given in Section A.2. The second line of Eq. 2.6 is the second-order RSD term. The contributions from local PNG or general relativity are omitted, apart from RSD – which is such an essential effect to account for that it is now generally included in the basic Newtonian framework.
@@
$\mathcal{K}_{\mathrm{nG}}^{(1)}(\bm{k}_{a})$ $=\frac{b_{01}}{\mathcal{M}_{a}}~{}\mbox{where}~{}\mathcal{M}_{a}\equiv\mathcal{M}(k_{a})=\sqrt{\frac{25\,P(k_{a})}{9\,P_{\cal R}(k_{a})}}=\frac{10\,D}{3\,\mathcal{H}^{2}\,(3\varOmega_{\rm m}+2f)}\,T(k_{a})\,k_{a}^{2}\;,$ (2.11)
@@ 3.1 Signal-to-noise ratio
$D_{P}(z,\bm{k}_{a})$ $=\exp\left\{-\frac{1}{2}\,\Big{[}k_{a}\,\mu_{a}\,\sigma(z)\Big{]}^{2}\right\}\;,$ (3.4)
$D_{B}(z,\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\exp\left\{-\frac{1}{2}\,\Big{[}k_{1}^{2}\,\mu_{1}^{2}+k_{2}^{2}\,\mu_{2}^{2}+k_{3}^{2}\,\mu_{3}^{2}\Big{]}\,\sigma(z)^{2}\right\}\;,$ (3.5)
@@
    $\left.\quad-\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{2})\,\mathcal{K}_{\rm N}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\right]\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{cp}\;.$ (3.7)
@@
    =$=\Big{\{}\mathcal{K}^{(1)}(\bm{k}_{1})\,\mathcal{K}^{(1)}(\bm{k}_{2})\,\mathcal{K}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$
@@
    $\quad+\left.\left(\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm GR}^{(1)}(\bm{k}_{2})\,+\mathcal{K}_{\rm GR}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{2})\right.\right.$
@@ 3.3 Parameter space
$\quad\left.+\epsilon^{(2)}_{\rm GR}\,GR^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,N^{(1)}(\bm{k}_{1},\bm{k}_{2})\right]\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{cp}\;,$ (3.11)
@@ 4.1 Galaxy survey specifications
Survey Redshift range Sky area [$\mathrm{deg}^{2}$]
Roman-like Model 1 & 3: $0.5\leq z\leq 2$ $2000$
DESI BGS $0\leq z\leq 0.7$ 14 0001400014\,00014 000
@@
15 0001500015\,00015 000
MegaMapper LBG $2.1\leq z\leq 5$ 20 0002000020\,00020 000
Table 1: Scope of chosen mock surveys
@@
$V(z)=\frac{4}{3}\,\pi\,f_{\rm sky}\,\left[\chi(z+\varDelta z/2)^{3}-\chi(z-\varDelta z/2)^{3}\right]\;,$ (4.1)
@@
Figure 4: Marginal errors (left: per redshift bin, right: cumulative) on the parameters $\epsilon_{\mathrm{GR}^{(1)}}$, $\epsilon_{\mathrm{GR}^{(2)}}$ and $f_{\mathrm{NL}}$ for $f_{\mathrm{NL}}=0$.
@@
Roman-like Model 1 0.285 0.569 3.970
Roman-like Model 3 0.251 0.494 4.308
DESI BGS 0.622 1.309 6.437
Euclid-like Model 1 0.133 0.242 2.610
Euclid-like Model 3 0.142 0.258 2.659
MegaMapper LBG 0.305 0.219 0.310
Table 2: Marginal errors on the parameters $\theta=\{\epsilon_{\rm GR}^{(1)},\epsilon_{\mathrm{GR}}^{(2)},f_{\mathrm{NL}}\}$ for each survey
@@
Implementing the parameter bias analysis of Section 3.4 for $-5<f_{\mathrm{NL}}<5$ yields Fig. 6. We show the biases for the total information matrix, i.e. the sum over all redshift bins.
@@
Figure 6: Total bias on parameter estimates for each survey, in units of marginal errors for $\epsilon_{\mathrm{GR}}^{(1)}$ (top), $\epsilon_{\mathrm{GR}}^{(2)}$ (centre) and $f_{\mathrm{NL}}$ (bottom) for $-5\leq f_{\mathrm{NL}}\leq 5$.
@@ 4.2 Including luminosity derivatives
Fig. 10 in Appendix D shows that including these terms also has an effect on the bias on $f_{\mathrm{NL}}$. Consequently, it is crucial to include the luminosity derivative terms for a complete analysis. This highlights the need for the investigation of the clustering bias dependence on luminosity for different galaxy targets.
@@ 5 Conclusion
The main conclusion that can be drawn from our analysis is that the higher the SNR of the GR signal, the more important it is to account for GR corrections when constraining $f_{\rm NL}$. This is shown most convincingly through Fig. 7. The least promising survey from our analysis, with the lowest precision for all three parameters, is the DESI BGS, which appears to be ineffective at detecting any GR corrections to the galaxy bispectrum within a reasonable margin of error. While this means that there is little bias on $f_{\rm NL}$ when neglecting GR effects, it also means a huge uncertainty in the measurement of $f_{\rm NL}$ compared to samples at greater depth – see the bottom left panel of Fig. 4.

### v3

@@ Abstract
Upcoming galaxy surveys aim to map the Universe with unprecedented precision, depth and sky coverage. The galaxy bispectrum is a prime source of information as it allows us to probe primordial non-Gaussianity (PNG), a key factor in differentiating various models of inflation. On the scales where local PNG is strongest, Doppler and other relativistic effects become important and need to be included. Unlike for the single-tracer power spectrum, the leading order imaginary Doppler term does not cancel out in the bispectrum, leaving a smoking gun imaginary dipole signal. We investigate the detectability and importance of relativistic and local PNG contributions in the galaxy bispectrum. We compute the signal-to-noise ratio for the detection of lightcone projection effects in the bispectrum . Furthermore, we perform information matrix forecasts on the local PNG parameter, $f_{\rm NL}$, and on the parametrised amplitudes of the relativistic corrections. Finally, we quantify the bias on the measurement of $f_{\rm NL}$ that arises from neglecting relativistic effects. Our results show that detections of both first- and second-order relativistic effects are promising with forthcoming spectroscopic survey specifications – and are largely unaffected by the uncertainty in $f_{\rm NL}$. Conversely, we show for the first time that neglecting relativistic corrections in the galaxy bispectrum can lead to a shift $1.5\,\sigma(f_{\rm NL})$ on the detected value of $f_{\rm NL}$, highlighting the importance of including relativistic effects in our modelling.
@@ 1 Introduction
Galaxy surveys have long been an indispensable probe of large-scale structure. We can infer the underlying physics that governs the evolution of matter perturbations by observing the galaxy distributions at different scales and redshifts, and therefore test the standard model of cosmology as well as theories of gravity. The anticipated measurements from current and future surveys such as the Dark Energy Spectroscopic Instrument (DESI) [1, 2], the Euclid satellite mission [3, 4], the Nancy Grace Roman Space Telescope [5, 6] and MegaMapper [7], will provide information on ultra-large scales, i.e. beyond the matter-radiation equality scale. At these scales primordial non-Gaussianity (PNG) can be probed. It is also at these scales that effects from general relativity have the most impact on our measurements. This is the key motivation of this paper, quantifying how corrections to the galaxy number density which appear due to general relativity, specifically those derived from observing on our past lightcone, affect measurements of PNG. We explore this through the galaxy bispectrum, the Fourier transform of the 3-point correlation function of galaxy density contrasts, which is only non-zero in the presence of non-Gaussianity. Doppler shifts induced by the peculiar velocities of galaxies, gravitational lensing effects, and other relativistic corrections alter the clustering patterns of galaxies, imprinting distinct signatures on the galaxy distribution which can mimic those of PNG [8, 9, 10, 11]. Therefore, understanding and accurately modeling these relativistic effects are imperative for extracting robust cosmological information from galaxy surveys.
@@
The galaxy power spectrum has been extensively studied and used to obtain impressive constraints on cosmological parameters. The dominant perturbation effect on sub-Hubble scales is from redshift-space distortions (RSD) [16], which constitute the standard Newtonian approximation to projection effects. The inclusion of relativistic corrections to the galaxy density contrast at linear order [17, 18, 19] has, in turn, provided further insights. Many studies have been done on the possibility of constraining the local PNG parameter $f_{\mathrm{NL}}$ using the relativistic galaxy power spectrum for future galaxy and intensity mapping surveys, which avoids the bias typical in the Newtonian analysis [8, 9, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. However, the leading order relativistic correction to the observed galaxy overdensity is the imaginary Doppler term, which, for a single-tracer, becomes real in the power spectrum, and is thus harder to differentiate from the PNG effects they mimic. Whereas, in the bispectrum, these terms induce a smoking gun imaginary dipole signal [36], making this higher order statistic a primary candidate for disentangling the two. PNG in the galaxy bispectrum has been extensively investigated in the Newtonian approximation [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. In [44] it was shown that constraints on $f_{\mathrm{NL}}$ can be obtained from large-scale structure which are competitive with that of Planck, using forecasts on the Newtonian tree-level bispectrum (including RSD to second order), based on survey specifications from BOSS [55], eBOSS [56], Euclid [57] and DESI (LRGs, ELGs and QSOs) [58].
@@ 2 The galaxy bispectrum
$\left<\varDelta_{\rm g}^{(1)}(\bm{k}_{1})\,\varDelta_{\rm g}^{(1)}(\bm{k}_{2})\,\varDelta_{\rm g}^{(2)}(\bm{k}_{3})\right>+2\,\mathrm{c.p.}=2\,(2\pi)^{3}\,B_{\rm g}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,\delta^{D}(\bm{k}_{1}+\bm{k}_{2}\,+\bm{k}_{3})\;,$ (2.1)
where $\mathrm{c.p.}$ denotes cyclic permutations over $\bm{k}_{1},\bm{k}_{2}$, and $\bm{k}_{3}$ and we use the convention that the observed density contrast is $\varDelta_{\rm g}=\varDelta_{\rm g}^{(1)}+\varDelta_{\rm g}^{(2)}/2$.
The bispectrum is related at first order to the number density contrast at the source $\delta_{\rm g}$ by
$\varDelta_{\rm g}=\delta_{\rm g}+\mathrm{RSD}+\text{relativistic effects}+\text{local PNG contributions}\;,$ (2.2)
and it can then be written in terms of first- and second-order Fourier kernels as
$B_{\rm g}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})=\mathcal{K}^{(1)}(\bm{k}_{1})\,\mathcal{K}^{(1)}(\bm{k}_{2})\,\mathcal{K}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{c.p.}\;,$ (2.3)
@@
$B_{\rm gN}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})=\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{2})\,\mathcal{K}_{\rm N}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{c.p.}\;,$ (2.4)
@@
Here, $\mu_{a}=\hat{\bm{k}}_{a}\cdot\hat{\bm{n}}$, with $\hat{\bm{n}}$ the line-of-sight direction, $f\coloneqq-{\rm d}\ln D/{\rm d}\ln(1+z)$ is the linear matter growth rate, and $D$ is the linear growth factor and $z$ the redshift. The kernels $F_{2}$ (mode-coupling part of the matter density contrast), $G_{2}$ (mode-coupling part of the matter velocity), and $S_{2}$ (tidal bias) are given in Section A.1, with the linear and quadratic Gaussian clustering biases $b_{10}$ and $b_{20}$ given in Section A.2. The second line of Eq. 2.6 is the second-order RSD term. The contributions from local PNG or general relativity are omitted, apart from RSD – which is such an essential effect to account for that it is now generally included in the basic Newtonian framework.
@@
$\mathcal{K}_{\mathrm{nG}}^{(1)}(\bm{k}_{a})$ $=\frac{b_{01}}{\mathcal{M}_{a}}~\mbox{where}~\mathcal{M}_{a}\equiv\mathcal{M}(k_{a}){\color[rgb]{0,0.88,0}=}\sqrt{\frac{25\,P(k_{a})}{9\,P_{\cal R}(k_{a})}}=\frac{10\,D}{3\,\mathcal{H}^{2}\,(3\,\varOmega_{\rm m}+2\,f)}\,T(k_{a})\,k_{a}^{2}\;,$ (2.11)
@@ 3.1 Signal-to-noise ratio
$D_{P}(z,\bm{k}_{a})$ $=\exp\left\{-\frac{1}{2}\,\left[k_{a}\,\mu_{a}\,\sigma(z)\right]^{2}\right\}\;,$ (3.4)
$D_{B}(z,\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\exp\left\{-\frac{1}{2}\,\left[k_{1}^{2}\,\mu_{1}^{2}+k_{2}^{2}\,\mu_{2}^{2}+k_{3}^{2}\,\mu_{3}^{2}\right]\,\sigma(z)^{2}\right\}\;,$ (3.5)
@@
    $\left.\quad-\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{1})\,\mathcal{K}_{\rm N}^{(1)}(\bm{k}_{2})\,\mathcal{K}_{\rm N}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\right]\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{c.p.}\;.$ (3.7)
@@
    $=\Big{\{}\mathcal{K}^{(1)}(\bm{k}_{1})\,\mathcal{K}^{(1)}(\bm{k}_{2})\,\mathcal{K}^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$
@@
[…]
@@ 3.3 Parameter space
$\quad\left.+\epsilon^{(2)}_{\rm GR}\,GR^{(2)}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\,N^{(1)}(\bm{k}_{1},\bm{k}_{2})\right]\,P(\bm{k}_{1})\,P(\bm{k}_{2})+2\,\mathrm{c.p.}\;,$ (3.11)
@@ 4.1 Galaxy survey specifications
Table 1: Specifications for the adopted mock surveys.
Survey Redshift range Sky fraction ($f_{\mathrm{sky}}$)
Roman-like Model 1 & 3: $5\leq z\leq 1.9 0.0$50.05
DESI BGS $0\leq z\leq 0.6$ $0.34$
@@
$0.36$
MegaMapper LBG $2.1\leq z\leq 5$ $0.50$
@@
$V(z)=\frac{4}{3}\,\pi\,f_{\rm sky}\,\left[\chi^{3}(z+\varDelta z/2)-\chi^{3}(z-\varDelta z/2)\right]\;,$ (4.1)
@@
Figure 4: Marginal errors (left: per redshift bin, right: cumulative) on the parameters $\epsilon_{\mathrm{GR}}^{(1)}$, $\epsilon_{\mathrm{GR}}^{(2)}$, and $f_{\mathrm{NL}}$ for $f_{\mathrm{NL}}=0$.
Table 2: Marginal errors on the parameters $\epsilon_{\rm GR}^{(1)},\epsilon_{\mathrm{GR}}^{(2)},f_{\mathrm{NL}}$ for each adopted survey.
@@
Roman-like Model 1 0.29 0.57 4.00
Roman-like Model 3 0.25 0.50 4.30
DESI BGS 0.81 1.60 6.40
Euclid-like Model 1 0.15 0.27 3.10
Euclid-like Model 3 0.14 0.26 2.70
MegaMapper LBG 0.32 0.21 0.26
@@
Implementing the parameter bias analysis of Section 3.4 for $5<f_{\mathrm{NL}}<5$ yields Fig. 6. We show the biases for the total information matrix, i.e. the sum over all redshift bins.
@@
Figure 6: Total bias on parameter estimates for each survey, in units of marginal errors for $\epsilon_{\mathrm{GR}}^{(1)}$ (top), $\epsilon_{\mathrm{GR}}^{(2)}$ (centre) and $f_{\mathrm{NL}}$ (bottom) for $5\leq f_{\mathrm{NL}}\leq 5$.
@@ 4.2 Including luminosity derivatives
Figure 10 in Appendix D shows that including these terms also has an effect on the bias on $f_{\mathrm{NL}}$. Consequently, it is crucial to include the luminosity derivative terms for a complete analysis. This highlights the need for the investigation of the clustering bias dependence on luminosity for different galaxy targets.
@@ 5 Conclusion
In this paper, we have investigated the detectability of relativistic corrections to the galaxy bispectrum in the presence of local PNG using analytically derived parameters of current and future galaxy surveys. The main conclusion that can be drawn from our analysis is that the higher the SNR of the GR signal, the more important it is to account for GR corrections when constraining $f_{\rm NL}$. This is shown most convincingly through Fig. 7. The least promising survey from our analysis, with the lowest precision for all three parameters, is the DESI BGS, which appears to be ineffective at detecting any GR corrections to the galaxy bispectrum within a reasonable margin of error. While this means that there is little bias on $f_{\rm NL}$ when neglecting GR effects, it also means a huge uncertainty in the measurement of $f_{\rm NL}$ compared to samples at greater depth – see the bottom left panel of Fig. 4.
@@
This work is the first to forecast constraints on relativistic amplitudes and $f_{\rm NL}$ using an analytical model of the observed galaxy bispectrum with local projection effects in the presence of local PNG. We aim to highlight the information and constraining power on $f_{\rm NL}$ which can be obtained uniquely through bispectrum measurements within the coming data releases from galaxy surveys, if relativistic effects are carefully modelled. A multi-tracer and a joint $P+B$ approach is left for future work.